OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle with legs of lengths 6 and 8. If necessary, round your answer to two decimal places. help Please

OpenStudy (anonymous):

We can find the length of the hypotenuse using the Pythagorean Formula: \[c = \sqrt{a^{2}+b^{2}} \] Plug in known values to get:\[c = \sqrt{6^{2}+8^{2}} \]\[c = \sqrt{36+64} \]\[c = \sqrt{100} \]\[c = 10\] The length of the hypotenuse is 10.

OpenStudy (anonymous):

Thank you